OpenStudy (candystarlover2):

Jeremy has 2 less than 3 times the number of scarves, x, Dawnie has. Which expression represents the total number of scarves Jeremy has? A. 2x−3 B. 3x−2 C. 3−2x D. 2−3x

OpenStudy (templeguy):

I'll answer yours if you answer mine

OpenStudy (candystarlover2):

whats yours? :)

OpenStudy (templeguy):

I'll send it in a message

OpenStudy (candystarlover2):

ok

OpenStudy (templeguy):

I sent it

OpenStudy (candystarlover2):

the link does not work :/

OpenStudy (candystarlover2):

Jeremy has 2 less than 3 times the number of scarves, x, Dawnie has. Which expression represents the total number of scarves Jeremy has? A. 2x−3 B. 3x−2 C. 3−2x D. 2−3x

OpenStudy (templeguy):

I resent the link

OpenStudy (candystarlover2):

I already found help at Brainly! the best ever!

OpenStudy (lowkey.s):

The answer is D. 2-3x. 2 less=2- 3 times the number of x=3x 2-3x

alones (alones):

@templeguy Don't send direct answers in pm's.

OpenStudy (lowkey.s):

Argeed @AloneS